If mama drives a car in a circle path, from point A and come back to the point A. With path's long is 1000 m and velocity is 100m/s, time's taken is 10 s. Hence, Determine the displacement vs time's graph


1
Expert's answer
2021-06-24T16:14:00-0400

"L=2R\cdot\sin(\frac{\alpha}{2})=2R\cdot\sin(\frac{\omega t}{2})=2R\cdot\sin(\frac{2\pi t}{2T})="


"=2R\cdot\sin(\frac{\pi t}{10})"


"l=2\pi R\to R=l/(2\pi)=1000/(2\pi)=159.2\ (m)" . So, the displacement vs time


"L=2\cdot 159.2\cdot\sin(\frac{\pi t}{10})=318.3\cdot\sin(\frac{\pi t}{10})" . Answer